A rectangular rug has the length of 12 feet and a diagonal of 20 feet. What is the perimeter of the rug?
Answer by venugopalramana(3286)

A rectangular rug has the length of 12 feet and a diagonal of 20 feet. What is the perimeter of the rug?
diagonal^2 = l^2+b^2
20^2=12^2+b^2
b^2=400-144=256
b=16'
perimeter = 2(l+b)=2(12+16)=56'